# On five questions of Bourin and Lee: symmetric moduli and an Euler operator identity

### v2

@@ 1.3. Matrix analogues of Euler’s quadrilateral identity
1.4. Sharp Clarkson–McCarthy type inequalities on Euler’s identity
For a Hilbert space $\mathcal{H}$, let $S_{p}$ denote the Schatten $p$-class. For $p>0$, we say that $X\in S_{p}$ if $X$ is a compact operator on $\mathcal{H}$ whose singular values $(\mu_{k}^{\downarrow}(X))_{k\geq 1}$ satisfy $\sum_{k\geq 1}\mu_{k}^{\downarrow}(X)^{p}<\infty$ . In this case, the Schatten pp-(quasi)norm is defined by
$\|X\|_{p}:=\Big(\sum_{k\geq 1}\mu_{k}^{\downarrow}(X)^{p}\Big)^{1/p}.$
(When $0<p<1$, $\|\cdot\|_{p}$ is a quasi-norm.)
Building on the Euler operator identity (1.6), we obtain the following sharp Clarkson–McCarthy type inequalities.
Theorem 1.19.
Let $A,B,C\in S_{p}$. Then for $p\geq 2$,
$\|A+B\|_{p}^{p}+\|B+C\|_{p}^{p}+\|C+A\|_{p}^{p}\leq 2^{p-2}\big(\|A+B+C\|_{p}^{p}+\|A\|_{p}^{p}+\|B\|_{p}^{p}+\|C\|_{p}^{p}\big).$
For $0<p\leq 2$, the inequality is reversed. Moreover, the constant $2^{p-2}$ is optimal: equality holds for all $p>0$ by taking $A=B=-C\neq 0$.
As a complement to Theorem 1.19, numerical evidence suggests the following sharp inequality. For a weaker bound with coefficient $3^{\frac{p}{2}-1}$, see Theorem 7.10.
Conjecture 1.20.
Let $A,B,C\in S_{p}$. Then for $p\geq 2$,
$\|A+B+C\|_{p}^{p}+\|A\|_{p}^{p}+\|B\|_{p}^{p}+\|C\|_{p}^{p}\leq\frac{3^{p-1}+1}{2^{p}}\left(\|A+B\|_{p}^{p}+\|B+C\|_{p}^{p}+\|C+A\|_{p}^{p}\right).$
For $0<p\leq 2$, the inequality is reversed. Moreover, the constant $\frac{3^{p-1}+1}{2^{p}}$ is optimal: equality holds for all $p>0$ by taking $A=B=C\neq 0$.
Classical Hlawka-type inequalities compare norms (or their powers) of vector sums with those of their pairwise sums. A standard form of Hlawka’s inequality [17] asserts that if $V$ is an inner product space and $x,y,z\in V$, then
$\|x+y\|+\|y+z\|+\|z+x\|\ \leq\ \|x\|+\|y\|+\|z\|+\|x+y+z\|,$
where $\|\cdot\|$ denotes the norm induced by the inner product; equivalently, the inequality holds for any norm satisfying the parallelogram identity. Numerous refinements and variants replace $\|\cdot\|$ by powers $\|\cdot\|^{p}$ and/or adjust the constants.
Theorem 1.19 and Conjecture 1.20 can be viewed as a Schatten–$p$ (i.e. noncommutative $L_{p}$) analogue: it features a sharp $p$-dependent constant and exhibits a change in behavior at $p=2$, as is typical for inequalities governed by the uniform convexity/concavity of $L_{p}$ spaces.
As a further refinement of Theorem 1.19, we establish the following sharp mixed $\ell_{q}$$\ell_{p}$ Clarkson–McCarthy type inequalities.
Theorem 1.21.
Let $A,B,C\in S_{p}$. Then for $1<p\leq 2$ ,
$\|A+B+C\|_{p}^{q}+\|A\|_{p}^{q}+\|B\|_{p}^{q}+\|C\|_{p}^{q}$ $\leq 2^{\,1-\frac{q}{p}}\,\Big(\|A+B\|_{p}^{p}+\|B+C\|_{p}^{p}+\|C+A\|_{p}^{p}\Big)^{q/p},$
$\|A+B\|_{p}^{q}+\|B+C\|_{p}^{q}+\|C+A\|_{p}^{q}$ $\leq 2^{\,1-\frac{q}{p}}\,\Big(\|A+B+C\|_{p}^{p}+\|A\|_{p}^{p}+\|B\|_{p}^{p}+\|C\|_{p}^{p}\Big)^{q/p},$
where $q$ is the conjugate exponent of $p$. For $p\geq 2$, these two inequalities are reversed. Moreover, the constant $2^{1-\frac{q}{p}}$ is optimal: equality holds for all $p>1$ by taking $A=B=-C\neq 0$.
@@ 1.4. Sharp Clarkson–McCarthy type inequalities on Euler’s identity
Setting C=BC=-B in Theorems 1.19 and 1.21 recovers the classical Clarkson–McCarthy inequality of [16]. For a complete characterization of the validity of Clarkson–McCarthy type inequalities, see the author’s paper [21, Corollary 1.14] and the following Theorem 1.23 (also see [21, Theorem 1.16]). For further results and developments on Clarkson–McCarthy type inequalities, we refer the reader to [5, 6, 9, 10, 15, 21, 22].
We now record the following mixed-norm inequality (Theorem 1.23), which strengthens the result in the author’s earlier work [21, Theorem 1.16]. Unlike the approach in [21], which applies the three-lines theorem to a suitable trace functional and then invokes a duality argument, we present a new proof based on a vector-valued Riesz–Thorin type complex interpolation scheme.
Theorem 1.23 (Mixed $\ell_{q}(S_{p})$$\ell_{p}(S_{p})$ estimates).
Let $U=[u_{ij}]\in\mathbb{M}_{s,t}$ and set
$\mu:=\max_{1\leq i\leq s,\ 1\leq j\leq t}|u_{ij}|.$
For $1<p<\infty$, let $q$ be the conjugate exponent of $p$, i.e. $\frac{1}{p}+\frac{1}{q}=1$. Define the linear map on $t$-tuples
$T_{U}:\ \ell_{p}^{t}(S_{p})\longrightarrow\ell_{q}^{s}(S_{p}),\qquad T_{U}(Z_{1},\dots,Z_{t}):=\Big(\sum_{j=1}^{t}u_{1j}Z_{j},\ \dots,\ \sum_{j=1}^{t}u_{sj}Z_{j}\Big).$
We will regard $T_{U}$ as an operator between $\ell_{p}$- and $\ell_{q}$-sums when equipped with the corresponding norms.
  1. (i)
    Assume that $U$ is a contraction, i.e. $\|U\|_{\infty}\leq 1$ (in particular, this holds if $U^{*}U=I_{t}$ or $UU^{*}=I_{s}$). Then for $1<p\leq 2$,
    $\Bigg(\sum_{i=1}^{s}\Big\|\sum_{j=1}^{t}u_{ij}Z_{j}\Big\|_{p}^{q}\Bigg)^{1/q}\ \leq\ \mu^{\,\frac{2}{p}-1}\,\Bigg(\sum_{j=1}^{t}\|Z_{j}\|_{p}^{p}\Bigg)^{1/p},$ (1.8)
    and the inequality is reversed for $p\geq 2$ if $U^{*}U=I_{t}$.
  2. (ii)
    Assume in addition that $U^{*}U=I_{t}$ (i.e. $U$ is an isometry). Then for $1<p\leq 2$,
    $\Bigg(\sum_{j=1}^{t}\|Z_{j}\|_{p}^{q}\Bigg)^{1/q}\ \leq\ \mu^{\,\frac{2}{p}-1}\,\Bigg(\sum_{i=1}^{s}\Big\|\sum_{j=1}^{t}u_{ij}Z_{j}\Big\|_{p}^{p}\Bigg)^{1/p},$ (1.9)
    and the inequality is reversed for $p\geq 2$.
Theorem 1.23 has several interesting consequences. For example, Audenaert–Kittaneh’s conjecture [1, Section 8.1, Conjecture 7] (proved by the author in [22]) states that if $A_{1},\ldots,A_{n}\in S_{p}$, then for $1<p\leq 2$,
$\left\|\sum_{i=1}^{n}A_{i}\right\|_{p}^{q}+\sum_{1\leq i<j\leq n}\|A_{i}-A_{j}\|_{p}^{q}\ \leq\ n\left(\sum_{i=1}^{n}\|A_{i}\|_{p}^{p}\right)^{q/p},$ (1.10)
where $q$ is the conjugate exponent of $p$; for $p\geq 2$ the inequality is reversed. Moreover, the constant $n$ is optimal: equality holds for all $1<p<\infty$ by taking $A_{1}=\ldots=A_{n}\neq 0$. This inequality can be recovered from Theorem 1.23 by choosing the isometry
$U=\frac{1}{\sqrt{n}}\begin{pmatrix}1&1&\cdots&1\\ 1&-1&\cdots&0\\ 1&0&\ddots&0\\ \vdots&\vdots&\ddots&\vdots\\ 0&0&\cdots&-1\end{pmatrix}\in\mathbb{M}_{\frac{n(n-1)+2}{2},\,n},$
whose rows encode the combinations $\sum_{i=1}^{n}A_{i}$ and $A_{i}-A_{j}$. Moreover, Theorem 1.23 yields the following sharp reverse-type companion to (1.10). For more details, see Subsection 7.4.
Corollary 1.24.
Let $A_{1},\ldots,A_{n}\in S_{p}$. Then for $1<p\leq 2$,
$\sum_{i=1}^{n}\|A_{i}\|^{q}_{p}\leq n^{-q/p}\left(\left\|\sum_{i=1}^{n}A_{i}\right\|_{p}^{p}+\sum_{1\leq i<j\leq n}\|A_{i}-A_{j}\|_{p}^{p}\right)^{\frac{q}{p}},$ (1.11)
where $q$ stands for the conjugate exponent of $p$. For $p\geq 2$, (1.11) is reversed. Moreover, the constant $n^{-q/p}$ is optimal: equality holds for all $1<p<\infty$ by taking $A_{1}=\ldots=A_{n}\neq 0$.
As further consequences of Euler’s identity (1.6), we next derive sharp Clarkson–McCarthy type inequalities for general unitarily invariant norms, together with a unitary-orbit refinement.
Theorem 1.25.
Let $A,B,C\in\mathbb{M}_{n}$. Then for $p\geq 2$,

### v3

@@ 1.4. Questions of Bourin and Lee on symmetric moduli and Euler-type identities
1.5. Future directions
After early versions of this paper were completed, some related works appeared. Bourin and Lee [11] established triangle inequalities for unitarily invariant norms on arithmetic symmetric modulus, while the author [21] studied triangle inequalities for unitarily invariant norms and Schatten $p$-norms on two symmetric moduli. Future directions for further study include the Lipschitz continuity of these moduli and related problems.
1.6. Organization of this paper
The paper is organized as follows. Section 2 answers Question 1.6 and proves Theorem 1.8. Section 3 establishes a rectangular Thompson’s inequality and uses it to prove Theorems 1.9 and 1.10. Section 4 extends the quadratic symmetric Thompson inequality to operators on Hilbert spaces and proves the corresponding equality characterization. Section 5 proves the concave-function subadditivity theorem for the quadratic symmetric modulus and derives its norm and trace consequences. Section 6 obtains condition-number estimates for the symmetric moduli and shows that the coefficient is sharp. Section 7 develops Clarkson–McCarthy type inequalities for the quadratic symmetric modulus, both in operator order and for unitarily invariant norms. Sections 8, 9, and 10 answer three questions raised in [10]. Finally, Section 11 establishes Euler-type operator identities together with several isometry-orbit refinements.
Acknowledgments
The author is very grateful to Professor Bourin for his valuable suggestions. This work is supported by the China Scholarship Council, the Young Elite Scientists Sponsorship Program for PhD Students (China Association for Science and Technology), and the Fundamental Research Funds for the Central Universities at Xi’an Jiaotong University (Grant No. xzy022024045).
2. A counterexample to Question 1.6 and proof of Theorem 1.8
In this section, we answer Question 1.6 and establish Theorem 1.8.
Answer to Question 1.6. The answer is negative. Indeed, consider
$f(t)=t-t^{2},\qquad t\geq 0.$
Then $f$ is concave on $0,\infty)$ and $f(0)=0$. Let
$A=\begin{pmatrix}1&0\\ 0&0\end{pmatrix},\qquad B=\begin{pmatrix}c^{2}&cs\\ cs&s^{2}\end{pmatrix},\qquad 0<c<1,\quad s=\sqrt{1-c^{2}}.$
Since $A$ and $B$ are rank-one projections, we have $A^{2}=A$ and $B^{2}=B$, and therefore
$f(A)=A-A^{2}=0,\qquad f(B)=B-B^{2}=0.$
Thus, for all unitaries $U,V$,
$Uf(A)U^{*}+Vf(B)V^{*}=0.$
If the conclusion of Theorem 1.5 were valid under the weaker assumption $f(0)=0$ and concavity alone, then we would have $f(A+B)\leq 0$.
However, $A+B$ has trace $2$ and determinant $1-c^{2}$, so its eigenvalues are $1\pm c$. Consequently, the eigenvalues of $f(A+B)$ are
$f(1+c)=-c(1+c)\qquad\text{and}\qquad f(1-c)=c(1-c).$
Since $0<c<1$, the latter is strictly positive. Hence $f(A+B)$ has a positive eigenvalue, so $f(A+B)\nleq 0$. Therefore there do not exist unitaries $U,V$ such that
$f(A+B)\leq Uf(A)U^{*}+Vf(B)V^{*}.$
This proves that Theorem 1.5 does not extend to arbitrary concave functions satisfying only $f(0)=0$. In particular, the monotonicity assumption cannot be removed.∎
Proof of Theorem 1.8.
Take
$A=\begin{pmatrix}-1&-1\\ 0&-1\end{pmatrix},\qquad B=\begin{pmatrix}0&-1\\ 0&0\end{pmatrix}\in\mathbb{M}_{2}.$
A direct computation gives
$|A|_{\mathrm{sym}}=\frac{1}{2\sqrt{5}}\begin{pmatrix}5&2\\ 2&5\end{pmatrix},\qquad|B|_{\mathrm{sym}}=\frac{1}{2}I_{2},\qquad|A+B|_{\mathrm{sym}}=\frac{1}{\sqrt{2}}\begin{pmatrix}2&1\\ 1&2\end{pmatrix}$.
Hence
$\bigl\||A+B|_{\mathrm{sym}}\bigr\|_{\infty}=\frac{3}{\sqrt{2}},\qquad\bigl\||A|_{\mathrm{sym}}\bigr\|_{\infty}=\frac{7}{2\sqrt{5}},\qquad\bigl\||B|_{\mathrm{sym}}\bigr\|_{\infty}=\frac{1}{2},$
so
$\bigl\||A+B|_{\mathrm{sym}}\bigr\|_{\infty}>\bigl\||A|_{\mathrm{sym}}\bigr\|_{\infty}+\bigl\||B|_{\mathrm{sym}}\bigr\|_{\infty}\qquad\left(\text{i.e. }\ \frac{3}{\sqrt{2}}>\frac{7}{2\sqrt{5}}+\frac{1}{2}\right).$
If the inequality in Theorem 1.8 held for some unitaries $U,V$, then taking operator norms would yield
$\bigl\||A+B|_{\mathrm{sym}}\bigr\|_{\infty}\leq\bigl\|U|A|_{\mathrm{sym}}U^{*}+V|B|_{\mathrm{sym}}V^{*}\bigr\|_{\infty}\leq\bigl\||A|_{\mathrm{sym}}\bigr\|_{\infty}+\bigl\||B|_{\mathrm{sym}}\bigr\|_{\infty},$
a contradiction. ∎
3. A rectangular Thompson’s inequality and proof of Theorem 1.9
In this section, we establish a rectangular version of Thompson’s inequality, which will be the main tool in the proof of Theorem 1.9. The key idea is to reduce the rectangular setting to the square one by means of the polar decomposition: for a rectangular matrix $X\in\mathbb{M}_{m,n}$, one can write $X=W|X|$ with a partial isometry $W$ whose initial projection coincides with the support projection of $|X|$. We then use the rectangular Thompson’s inequality to prove Theorem 1.9.
Lemma 3.1.
Let $A\in\mathbb{M}_{m,n}$. Then there exists a partial isometry $W\in\mathbb{M}_{m,n}$ such that
$A=W|A|,\qquad|A|=(A^{*}A)^{1/2},$
and
$W^{*}W=\mathrm{supp}(|A|),\qquad W^{*}W\leq I_{n},WW^{*}\leq I_{m},\qquad(W^{*}W)\,|A|=|A|,$
where $\mathrm{supp}(\cdot)$ means the support projection.
Proof.
Take a singular value decomposition $A=U\Sigma V^{*}$, where $U\in\mathbb{M}_{m}$ and $V\in\mathbb{M}_{n}$ are unitary and
$\Sigma=\begin{pmatrix}D&0\\ 0&0\end{pmatrix}\in\mathbb{M}_{m,n},\qquad D=\operatorname{diag}(s_{1},\dots,s_{r}),\ \ s_{1}\geq\cdots\geq s_{r}>0,\ \ r=\operatorname{rank}(A).$
Then
$A^{*}A=V(\Sigma^{*}\Sigma)V^{*},\qquad\Sigma^{*}\Sigma=\begin{pmatrix}D^{2}&0\\ 0&0\end{pmatrix}$,
hence
$|A|=(A^{*}A)^{1/2}=V\begin{pmatrix}D&0\\ 0&0\end{pmatrix}V^{*}.$
$W|A|=U\begin{pmatrix}I_{r}&0\\ 0&0\end{pmatrix}V^{*}\cdot V\begin{pmatrix}D&0\\ 0&0\end{pmatrix}V^{*}=U\begin{pmatrix}D&0\\ 0&0\end{pmatrix}V^{*}=U\Sigma V^{*}=A.$
Moreover,
$(W^{*}W)\,|A|=\Bigl(V\begin{pmatrix}I_{r}&0\\ 0&0\end{pmatrix}V^{*}\Bigr)\Bigl(V\begin{pmatrix}D&0\\ 0&0\end{pmatrix}V^{*}\Bigr)=V\begin{pmatrix}D&0\\ 0&0\end{pmatrix}V^{*}=|A|.$
That is, $W^{*}W$ is exactly the support projection of $|A|$. ∎
Lemma 3.1 allows us to choose the partial isometry in the polar decomposition so that it acts as the identity on the range of $|A|$. This normalization is exactly what we need to turn $|A+B|$ into the absolute value of a square matrix, to which the classical Thompson inequality applies.
Proposition 3.2.
Let $A,B\in\mathbb{M}_{m,n}$. Then there exist two unitaries $U,V$ such that
$|A+B|\leq U|A|U^{*}+V|B|V^{*}.$
Proof.
$W^{*}(A+B)=W^{*}\left(W|A+B|\right)=|A+B|.$
In particular,
$|A+B|=|W^{*}(A+B)|=|W^{*}A+W^{*}B|.$ (3.1)
Now $W^{*}A$ and $W^{*}B$ are $n\times n$ (square) matrices. Applying Thompson’s inequality (Theorem 1.1) to the square matrices $W^{*}A$ and $W^{*}B$, there exist unitaries $U,V\in\mathbb{M}_{n}$ such that
$|W^{*}A+W^{*}B|\leq U\,|W^{*}A|\,U^{*}+V\,|W^{*}B|\,V^{*}.$ (3.2)
Since $WW^{*}\leq I_{m}$, we have
$|W^{*}A|^{2}=(W^{*}A)^{*}(W^{*}A)=A^{*}WW^{*}A\leq A^{*}A=|A|^{2}.$
$U|W^{*}A|U^{*}\leq U|A|U^{*},\qquad V|W^{*}B|V^{*}\leq V|B|V^{*}.$
Substituting into (3.2) and combining with (3.1) yields our desired result. ∎
We now prove Theorem 1.9, it follows from an application of Proposition 3.2 to a suitable linear embedding that converts the quadratic symmetrization into a rectangular modulus.
Proof of Theorem 1.9.
Define the linear map $\Gamma:\mathbb{M}_{n}\to\mathbb{M}_{2n,n}$ by
$\Gamma(T):=\frac{1}{\sqrt{2}}\begin{pmatrix}T\\ T^{*}\end{pmatrix}$.
Then
$\Gamma(A+B)=\Gamma(A)+\Gamma(B)\quad\text{ and }\quad|\Gamma(T)|=\,|T|_{\mathrm{qsym}}.$
Applying rectangular Thompson’s inequality (Proposition 3.2) yields unitaries $U,V\in\mathbb{M}_{n}$ such that
$|\Gamma(A+B)|$ $=|\Gamma(A)+\Gamma(B)|$
$\leq U|\Gamma(A)|U^{*}+V|\Gamma(B)|V^{*}.$
That is,
$|A+B|_{\mathrm{qsym}}\leq U|A|_{\mathrm{qsym}}U^{*}+V|B|_{\mathrm{qsym}}V^{*}.$
We now turn to the equality case. The argument is most transparent after embedding the rectangular problem into a square one, where Thompson’s equality theorem is available.
Proof of Theorem 1.10.
Define
$A_{0}:=\Gamma(A)\in\mathbb{M}_{2n,n},\qquad B_{0}:=\Gamma(B)\in\mathbb{M}_{2n,n}.$
Then
$|A_{0}|=\left|A\right|_{\mathrm{qsym}},\qquad|B_{0}|=\left|B\right|_{\mathrm{qsym}},\qquad|A_{0}+B_{0}|=\left|A+B\right|_{\mathrm{qsym}}.$
Hence the assumption
$\left|A+B\right|_{\mathrm{qsym}}=U\left|A\right|_{\mathrm{qsym}}U^{*}+V\left|B\right|_{\mathrm{qsym}}V^{*}$
is equivalent to
$|A_{0}+B_{0}|=U|A_{0}|U^{*}+V|B_{0}|V^{*}.$ (3.3)
Step 1: Embed the rectangular problem into a square one. Consider the block embedding of $\mathbb{M}_{2n,n}$ into $\mathbb{M}_{3n}$ given by
$\widetilde{A}:=\begin{pmatrix}0&0\\ A_{0}&0\end{pmatrix},\qquad\widetilde{B}:=\begin{pmatrix}0&0\\ B_{0}&0\end{pmatrix}\quad\text{in }\mathbb{M}_{3n},$
with respect to the decomposition $\mathbb{C}^{3n}\cong\mathbb{C}^{n}\oplus\mathbb{C}^{2n}$. A direct computation shows
$|\widetilde{A}|=\operatorname{diag}(|A_{0}|,0),\qquad|\widetilde{B}|=\operatorname{diag}(|B_{0}|,0),\qquad|\widetilde{A}+\widetilde{B}|=\operatorname{diag}(|A_{0}+B_{0}|,0).$
Let
$\widetilde{U}:=\operatorname{diag}(U,I_{2n}),\qquad\widetilde{V}:=\operatorname{diag}(V,I_{2n})\quad\text{in }\mathbb{M}_{3n}.$
Then (3.3) implies
$|\widetilde{A}+\widetilde{B}|=\operatorname{diag}(|A_{0}+B_{0}|,0)=\operatorname{diag}(U|A_{0}|U^{*}+V|B_{0}|V^{*},0)=\widetilde{U}|\widetilde{A}|\widetilde{U}^{*}+\widetilde{V}|\widetilde{B}|\widetilde{V}^{*}$.
Step 2: Apply Thompson’s equality theorem in $\mathbb{M}_{3n}$. By Theorem 1.2, there exists a unitary $\widetilde{W}\in\mathbb{M}_{3n}$ such that
[…]
Write $\widetilde{W}$ in block form (with respect to $\mathbb{C}^{n}\oplus\mathbb{C}^{2n}$):
$\widetilde{W}=\begin{pmatrix}W_{11}&W_{12}\\ W_{21}&W_{22}\end{pmatrix}$.
[…]
$\begin{pmatrix}0&0\\ A_{0}&0\end{pmatrix}=\begin{pmatrix}W_{11}&W_{12}\\ W_{21}&W_{22}\end{pmatrix}\begin{pmatrix}|A_{0}|&0\\ 0&0\end{pmatrix}=\begin{pmatrix}W_{11}|A_{0}|&0\\ W_{21}|A_{0}|&0\end{pmatrix}$.
Hence
$A_{0}=W_{21}|A_{0}|\qquad\text{and}\qquad W_{11}|A_{0}|=0.$
[…]
$B_{0}=W_{21}|B_{0}|\qquad\text{and}\qquad W_{11}|B_{0}|=0.$
Step 3: Construct the common partial isometry. Let
$P:=\operatorname{supp}(|A_{0}|+|B_{0}|)\in\mathbb{M}_{n}.$
Then $P|A_{0}|=|A_{0}|$ and $P|B_{0}|=|B_{0}|$, so if we define
$W:=W_{21}P\in\mathbb{M}_{2n,n},$
then
$W|A_{0}|=W_{21}P|A_{0}|=W_{21}|A_{0}|=A_{0},\qquad W|B_{0}|=W_{21}P|B_{0}|=W_{21}|B_{0}|=B_{0}.$
Moreover, since $W_{11}|A_{0}|=0$ and $W_{11}|B_{0}|=0$, we have
$W_{11}(|A_{0}|+|B_{0}|)=0,$
hence $W_{11}P=0$. Therefore
$\widetilde{W}\,\operatorname{diag}(P,0)=\begin{pmatrix}W_{11}P&0\\ W_{21}P&0\end{pmatrix}=\begin{pmatrix}0&0\\ W&0\end{pmatrix}$.
Because $\widetilde{W}$ is unitary, $\widetilde{W}\,\operatorname{diag}(P,0)$ is a partial isometry with initial projection $\operatorname{diag}(P,0)$. Taking adjoints and multiplying yields
$\begin{pmatrix}0&W^{*}\\ 0&0\end{pmatrix}\begin{pmatrix}0&0\\ W&0\end{pmatrix}=\operatorname{diag}(P,0),$
and hence
$W^{*}W=P.$
So $W$ is a partial isometry in $\mathbb{M}_{2n,n}$.
Finally, since $A_{0}=\Gamma(A)$, $B_{0}=\Gamma(B)$, $|A_{0}|=\left|A\right|_{\mathrm{qsym}}$, and $|B_{0}|=\left|B\right|_{\mathrm{qsym}}$, we conclude that
$\Gamma(A)=W\,\left|A\right|_{\mathrm{qsym}},\qquad\Gamma(B)=W\,\left|B\right|_{\mathrm{qsym}}.$
Equivalently, the polar decompositions of $\Gamma(A)$ and $\Gamma(B)$ share the same partial isometry factor $W$. ∎
4. Infinite-dimensional extensions and equality cases
In this section, we discuss the infinite-dimensional case of the triangle inequality for the quadratic symmetric modulus.
We begin by recording a rectangular version of Theorem 1.3 for operators in $\mathcal{B}(\mathcal{H},\mathcal{K})$.
Proposition 4.1.
Let $\mathcal{H},\mathcal{K}$ be complex Hilbert spaces and let $A,B\in\mathcal{B}(\mathcal{H},\mathcal{K})$. Then there exist isometries $U,V\in\mathcal{B}(\mathcal{H})$ such that
$\left|A+B\right|\ \leq\ U\,\left|A\right|\,U^{*}\;+\;V\,\left|B\right|\,V^{*}.$
Proof.
Let $A+B=W\,\left|A+B\right|$ be the polar decomposition, where $W\in\mathcal{B}(\mathcal{H},\mathcal{K})$ is a partial isometry. Recall that for an operator $S$ we write $\operatorname{Ran}(S)$ for its range. For a positive operator $S$ we denote by $\operatorname{supp}(S)$ its support projection, i.e. the orthogonal projection onto $\overline{\operatorname{Ran}(S)}$. Then
$W^{*}(A+B)=W^{*}W\,\left|A+B\right|=\left|A+B\right|,$
since $W^{*}W=\operatorname{supp}(\left|A+B\right|)$. Hence
$\left|A+B\right|=\left|W^{*}(A+B)\right|=\left|W^{*}A+W^{*}B\right|.$
Now $W^{*}A,W^{*}B\in\mathcal{B}(\mathcal{H})$ are operators on $\mathcal{H}$, so by Akemann–Anderson–Pedersen’s inequality (Theorem 1.3) there exist isometries $U,V\in\mathcal{B}(\mathcal{H})$ such that
$\left|W^{*}A+W^{*}B\right|\ \leq\ U\,\left|W^{*}A\right|\,U^{*}\;+\;V\,\left|W^{*}B\right|\,V^{*}.$
Finally, since $WW^{*}\leq I_{\mathcal{K}}$ we have
$\left|W^{*}A\right|^{2}=(W^{*}A)^{*}(W^{*}A)=A^{*}WW^{*}A\leq A^{*}A=\left|A\right|^{2},$
hence $\left|W^{*}A\right|\leq\left|A\right|$ by operator monotonicity of $t\mapsto t^{1/2}$. Similarly $\left|W^{*}B\right|\leq\left|B\right|$. Conjugation preserves order, so $U\left|W^{*}A\right|U^{*}\leq U\left|A\right|U^{*}$ and $V\left|W^{*}B\right|V^{*}\leq V\left|B\right|V^{*}$. Combining the last three displays yields the desired inequality. ∎
Now, we give a proof of Theorem 1.11.
Proof of Theorem 1.11.
Define the linear lifting
$\Gamma:\mathcal{B}(\mathcal{H})\to\mathcal{B}(\mathcal{H},\mathcal{H}\oplus\mathcal{H}),\qquad\Gamma(T):=\frac{1}{\sqrt{2}}\binom{T}{T^{*}}$.
Then
$\Gamma(A+B)=\Gamma(A)+\Gamma(B)\quad\text{ and }\quad|\Gamma(T)|=\,|T|_{\mathrm{qsym}}.$
Applying Proposition 4.1 yields isometries $U,V\in\mathcal{B}(\mathcal{H})$ such that
$|\Gamma(A+B)|$ $=|\Gamma(A)+\Gamma(B)|$
$\leq U|\Gamma(A)|U^{*}+V|\Gamma(B)|V^{*}.$
That is,
$|A+B|_{\mathrm{qsym}}\leq U|A|_{\mathrm{qsym}}U^{*}+V|B|_{\mathrm{qsym}}V^{*}.$
@@ 4. Infinite-dimensional extensions and equality cases
In finite dimension, every isometry is unitary. Hence Theorem 1.11 implies Theorem 1.9.
Before proving Theorem 1.12, we need the standard two-space extension of Theorem 1.4 to operators between Hilbert spaces.
Proposition 4.3.
Let $\mathcal{H},\mathcal{K}$ be complex Hilbert spaces and let $A,B\in\mathcal{B}(\mathcal{H},\mathcal{K})$. If
$\left|A+B\right|=\left|A\right|+\left|B\right|,$
then there exists a partial isometry $W\in\mathcal{B}(\mathcal{H},\mathcal{K})$ such that
$A=W\left|A\right|,\qquad B=W\left|B\right|.$
Proof.
Consider $\mathcal{H}\oplus\mathcal{K}$ and define block operators
$\widetilde{A}:=\begin{pmatrix}0&0\\ A&0\end{pmatrix},\qquad\widetilde{B}:=\begin{pmatrix}0&0\\ B&0\end{pmatrix}\quad\text{in }\mathcal{B}(\mathcal{H}\oplus\mathcal{K}).$